# UVE: Are MLLMs Unified Evaluators for AI-Generated Videos?

### v2

@@ 5.3 Effect of Prompting Strategy
The results in Figure 3 (a) demonstrate that using simplified prompts (without aspect-specific descriptions) generally led to decreased performance in InternVL-2.5-78B-MPO. The performance decline was particularly pronounced in several categories: Temporal Quality subaspects, Structural Correctness, Technical Quality, and Light & Color. However, other aspects maintain relatively stable performance levels. This phenomenon suggests that it is generally beneficial to provide a detailed description of the aspect to evaluate. Results of other models also showcase an advantage of detailed aspect description on the average performance (Appendix B.1).
@@
Similarly, Figure 3 (b) shows that simplified prompts result in reduced performance for video pair comparisons, reinforcing the value of detailed aspect descriptions in the prompts. Interestingly, contrary to established findings in image quality comparison (Zhang et al., 2024f), removing video order indicators does not significantly impact performance. We attribute this robustness to the enhanced temporal order understanding capabilities of recent MLLMs.
5.4 Effect of Scoring Strategy
@@
The results of Qwen2-VL-7B, presented in Figure 3, reveal that: (1) good/bad demonstrates superior performance for Temporal Quality and Static Quality assessments, while yes/no performs better for Dynamic Degree evaluation. (2) The adaptive strategy yields the best overall results for Qwen2-VL-7B. However, for other advanced MLLMs like InternVL-2.5-78B-MPO, the improvement from adaptive scoring is minimal (see Appendix B.2). (3) Directly prompting MLLMs to generate rating scores performs significantly worse than probability-based scoring methods. These findings suggest that using yes/no as unified scoring tokens strikes a good balance between simplicity and effectiveness.
5.5 Effect of Frame Number
Figure 5 illustrates the impact of varying video frame numbers on model performance. We can see that: (1) In the zero-frame scenario (pure-text input), MLLMs perform poorly as their outputs remain constant regardless of visual information, confirming that language shortcuts cannot be exploited to achieve good performance in UVE-Bench. (2) As the frame count increases, aspects requiring temporal understanding (Light & Color) show more consistent improvement compared to aspects focusing on individual frame information (Aesthetic Quality). (3) The 72B-scale models demonstrate more consistent improvement in average accuracy with increasing frame numbers compared to their 7B-scale counterparts. This suggests that larger models possess superior capabilities in processing and integrating information from additional video frames.
5.6 Adapting Single Ratings for Pairwise Comparison
@@
While our previous experiments involve direct pairwise video comparisons by simultaneously feeding two videos into MLLMs, we now explore an alternative approach. This method first independently rates each video using Eq. 1 and then converts these individual ratings into a four-way selection from set $\mathbf{O}$. The detailed conversion methodology is discussed in Appendix A.3.
@@
As shown in Figure 5, adaptation from single video rating brings substantial improvement to Qwen2-VL-7B compared to directly performing video pair comparison. However, Qwen2-VL-72B shows only minimal improvements under the same approach. Similar patterns were observed with InternVL-2.5-MPO, as shown in Appendix B.3. We hypothesize that smaller-scale models, like the 7B variant, have inherently weaker pairwise comparison capabilities, and adapting from single video rating bypasses this limitation. Based on these findings, we recommend implementing the single-to-pairwise adaptation strategy particularly for MLLMs with limited pairwise comparison capabilities.
@@ 6 Conclusion and Future Work
This study investigates the potential of MLLMs to serve as unified evaluators for AI-generated videos. To answer this question, we introduce the UVE-Bench, a benchmark designed to assess the capability of unified AIGV evaluation. Compared to existing AIGV datasets, UVE-Bench highlights (1) comprehensive evaluation aspects, (2) videos produced by SOTA video generative models and (3) reliable evaluation of both single video rating and video pair comparison. Based on UVE-Bench, we extensively evaluates 16 MLLMs. Our findings reveal that while SOTA MLLMs cannot fully replace human evaluators, they significantly outperform existing specialized evaluation methods, showcasing promising potential as unified AIGV evaluators. Our analytical studies highlight several design choices, including: providing detailed aspect-specific description in the prompts (§5.3), using yes/no as the scoring tokens for single video rating (§5.4) and adapting single video rating for pairwise comparison when using 7B-scale MLLMs (§5.6).
Future work of this study will extend to additional evaluation scenarios, including safety, ethics, and bias assessment, as well as diverse video generation conditions such as image-to-video and video-to-video generation.
Impact Statement
@@
The evaluation of video generative models (VGMs) currently relies predominantly on human assessment (Polyak et al., 2024; Kong et al., 2024) or qualitative analysis (Brooks et al., 2024). Our research demonstrates that multimodal large language models (MLLMs) could potentially serve as unified evaluators for AI-generated videos. The transition from human to automated evaluation frameworks would have significant dual implications: On the one hand, this shift would substantially reduce evaluation costs, thereby accelerating the development cycle of VGMs. On the other hand, if VGMs are optimized solely based on model feedback, there is a potential risk of divergence between the automated evaluators’ preferences and human values. Therefore, it essential to carefully scrutinize the correlation between automatic metrics and humans in the future research on AIGV evaluation.
Appendix A More Details of Experimental Settings
A.1 Evaluated Systems
@@
We conduct zero-shot evaluations on 16 MLLMs, the detailed information of which is summarized in Table 6.
VideoScore-v1.1.
(He et al., 2024) This model is a fine-tuned version of Mantis-Idefics2-8B (Jiang et al., 2024), trained to imitate human rating scores. In their work, He et al. (2024) gathered human ratings (on a discrete scale from 1 to 4) for five key evaluation aspects: Visual Quality, Temporal Consistency, Dynamic Degree, Text-to-Video Alignment, and Factual Consistency. The VideoScore model is trained using a regression approach to predict these human rating scores. According to the definition of these aspects, we associate them with 8 out of the 15 aspects in UVE-Bench, as detailed in Table 7.
VBench.
(Huang et al., 2024) VBench is a benchmark suite that comprehensively evaluates 16 aspects of AIGV using off-the-shelf models and specifically designed rules. As shown in Table 7, we identified an overlap of 6 aspects between VBench and our UVE-Bench, and we adopted the VBench metrics to evaluate these 6 aspects in UVE-Bench.
A.2 Evaluation Criteria for Single Video Rating
@@
$\mathcal{A}^{\text{single}}=\begin{cases}\mathbf{1}(\mathcal{S}_{1}>\mathcal{S}_{2})&\text{if }\mathcal{P}=``\mathcal{V}_{1}\text{better}"\\ \mathbf{1}(\mathcal{S}_{1}<\mathcal{S}_{2})&\text{elif }\mathcal{P}=``\mathcal{V}_{2}\text{better}"\\ f_{c}(\mathcal{S}_{1}|\beta)\cdot f_{c}(\mathcal{S}_{2}|\beta)&\text{elif }\mathcal{P}=``\text{same good}"\\ f^{{}^{\prime}}_{c}(\mathcal{S}_{1}|\alpha)\cdot f_{c}(\mathcal{S}_{2}|\alpha)&\text{elif }\mathcal{P}=``\text{same bad}"\\ \end{cases}$ (4)
@@
where $s$ is the coefficient that controls the speed of decaying. The graph of $f_{c}(\mathcal{S}|\beta)$ and $f^{\prime}_{c}(\mathcal{S}|\alpha)$ are illustrated in Figure 6.
A.3 Methodology for Adapting Single Video Rating for Pairwise Comparison
@@
[…]
Appendix B More Experimental Results
B.1 Effect of Prompting Strategy
@@
Figure B.1 presents the results of different prompting strategy with InternVL-2.5-8B-MPO, Qwen2-VL-7B and Qwen2-VL-72B. Unlike InternVL-2.5-78B-MPO (previously presented in Figure 3), these three models showed a less pronounced advantage when using detailed aspect-specific descriptions, with the simplified prompt achieving better performance in several aspects. We conjecture that this is because these three models cannot comprehend the detailed description as effectively as InternVL-2.5-78B-MPO. Nevertheless, when considering overall performance, the full prompt consistently maintained a slight improvement over the simplified version across all three models. Consistent with our findings for InternVL-2.5-78B-MPO, removing video order indicator does not result in significant performance change.
B.2 Effect of Scoring Strategy
@@
In §5.4, we analyze the effect of different scoring strategies with Qwen2-VL-7B. Here, we extended the investigation to three additional models: Qwen2-VL-72B, InternVL-2.5-8B-MPO, and InternVL-2.5-78B. The results are presented in Figure 8. Our findings reveal consistent patterns across all models. Similar to Qwen2-VL-7B, the good/bad and yes/no scoring strategies demonstrate distinct strengths in different evaluation aspects. Direct score generation through MLLMs consistently yields suboptimal results across all models. However, compared with Qwen2-VL-7B, the advantage of adaptive scoring token is less pronounced in Qwen2-VL-72B, InternVL-2.5-8B-MPO and InternVL-2.5-78B. This suggests that these three models are more robust to the variation of scoring tokens. These findings confirm that using yes/no as unified scoring token is simple yet effective.
B.3 Adapting Single Ratings for Pairwise Comparison

### v4

@@ 4.3.1 Impact of Prompting Strategy
The results in Fig. 4(b) (a) demonstrate that using simplified prompts (without aspect-specific descriptions) generally led to decreased performance in InternVL-2.5-78B-MPO. The performance decline was particularly pronounced in several categories: Temporal Quality subaspects, Structural Correctness, Technical Quality, and Light & Color. However, other aspects maintain relatively stable performance levels. This phenomenon suggests that it is generally beneficial to provide a detailed description of the aspect to evaluate. Results of other models also showcase an advantage of detailed aspect description on the average performance (Appendix D.1).
(c) Results with varying numbers of video frames.
(d) Results of direct video pair comparison versus adapting single video rating to pairwise comparison.
@@
Similarly, Fig. 4(b) (b) shows that simplified prompts result in reduced performance for video pair comparisons, reinforcing the value of detailed aspect descriptions in the prompts. Interestingly, contrary to established findings in image quality comparison [77], removing video order indicators does not significantly impact performance. We attribute this robustness to the enhanced temporal order understanding capabilities of recent MLLMs.
4.3.2 Impact of Scoring Strategy
@@
The results of Qwen2-VL-7B, presented in Fig. 4(b), reveal that: (1) good/bad demonstrates superior performance for Temporal Quality and Static Quality assessments, while yes/no performs better for Dynamic Degree evaluation. (2) The adaptive strategy yields the best overall results for Qwen2-VL-7B. However, for other advanced MLLMs like InternVL-2.5-78B-MPO, the improvement from adaptive scoring is minimal (see Appendix D.2). (3) Directly prompting MLLMs to generate rating scores performs significantly worse than probability-based scoring methods. These findings suggest that using yes/no as unified scoring tokens strikes a good balance between simplicity and effectiveness.
4.3.3 Adapting Single Ratings for Pairwise Comparison
@@
While our previous experiments involve direct pairwise video comparisons by simultaneously feeding two videos into MLLMs, we now explore an alternative approach. This method first independently rates each video using Eq. 1 and then converts these individual ratings into a four-way selection from set $\mathbf{O}$. The detailed conversion methodology is discussed in Appendix C.3.
@@
As shown in Fig. 4(d), adaptation from single video rating brings substantial improvement to Qwen2-VL-7B compared to directly performing video pair comparison. However, Qwen2-VL-72B shows only minimal improvements under the same approach. Similar patterns were observed with InternVL-2.5-MPO, as shown in Appendix D.4. We hypothesize that smaller-scale models, like the 7B variant, have inherently weaker pairwise comparison capabilities, and adapting from single video rating bypasses this limitation. Based on these findings, we recommend implementing the single-to-pairwise adaptation strategy particularly for MLLMs with limited pairwise comparison capabilities.
4.3.4 Impact of Frame Number
Fig. 4(d) illustrates the impact of varying video frame numbers on model performance. We can see that: (1) In the zero-frame scenario (pure-text input), MLLMs perform poorly as their outputs remain constant regardless of visual information, confirming that language shortcuts cannot be exploited to achieve good performance in UVE-Bench. (2) As the frame count increases, aspects requiring temporal understanding (Light & Color) show more consistent improvement compared to aspects focusing on individual frame information (Aesthetic Quality). (3) The 72B-scale models demonstrate more consistent improvement in average accuracy with increasing frame numbers compared to their 7B-scale counterparts. This suggests that larger models possess superior capabilities in processing and integrating information from additional video frames.
Table 6: Automatic and human rating of video generative models. The human prefer rate is computed as the ratio of “better” plus “same good” examples, based on video pair preference annotation.
Open-Sora 1.2 CogVideoX-5B Movie Gen Video HunyuanVideo Real Videos Static Quality Auto (Qwen2-VL-72B) 63.2 70.0 84.3 84.9 85.7 Human (Prefer Rate) 3.3 21.6 96.4 94.7 - Temporal Quality Auto (Qwen2-VL-72B) 59.7 67.3 74.7 77.1 78.0 Human (Prefer Rate) 1.5 42.4 96.6 86.6 -
4.4 Evaluation of Real versus Generated Videos
To examine the difference when applying MLLMs to evaluate real videos instead of AI-generated videos, we collect 204 real videos from the TempCompass [41] dataset and utilize Qwen2-VL-72B to perform single video rating on two aspects: Static Quality and Temporal Quality. For comparison, we also evaluate videos generated by Open-Sora 1.2, CogVideoX-5B, Movie Gen Video and HunyuanVideo. As shown in Tab. 6, real videos consistently receive higher scores than generated ones across both quality aspects. Furthermore, the relative ranking of VGMs basically aligns with human perceptual judgments: the more recent models (Movie Gen Video and HunyuanVideo) outperform earlier ones (Open-Sora 1.2 and CogVideoX-5B) in both human and automatic evaluations. However, Qwen2-VL-72B and humans disagree slightly on the rating of Movie Gen Video and HunyuanVideo, highlighting the need for more advanced evaluators capable of finer-grained comparisons.
@@ 5 Conclusion and Future Work
This study investigates the potential of MLLMs to serve as unified evaluators for AI-generated videos. To answer this question, we introduce the UVE-Bench, a benchmark designed to assess the capability of unified AIGV evaluation. Compared to existing AIGV datasets, UVE-Bench highlights (1) comprehensive evaluation aspects, (2) videos produced by SOTA video generative models and (3) reliable evaluation of both single video rating and video pair comparison. Based on UVE-Bench, we extensively evaluates 18 MLLMs. Our findings reveal that while SOTA MLLMs cannot fully replace human evaluators, they significantly outperform existing specialized evaluation methods, showcasing promising potential as unified AIGV evaluators. Our analytical studies highlight several design choices, including: providing detailed aspect-specific description in the prompts (§4.3.1), using yes/no as the scoring tokens for single video rating (§4.3.2) and adapting single video rating for pairwise comparison when using 7B-scale MLLMs (§4.3.3).
Looking ahead, we identify two main directions for future work: First, we aim to extend the benchmark to cover broader AIGV evaluation scenarios, including safety, ethics, and bias assessments, as well as more diverse generation settings such as image-to-video and video-to-video tasks. Second, we plan to enhance MLLM performance in AIGV evaluation through three complementary directions: (1) Advancing video encoding techniques to capture richer temporal information without compromising efficiency, which is essential to the temporally sensitive evaluation aspects. (2) Automatically curating high- and low-quality video pairs to train evaluators. This can be achieved by leveraging VGMs at different training stages or intentionally varying prompts or other hyperparameters during inference to generate controlled degradations in videos. (3) Incorporating visual generative loss into MLLM training—alongside language generation loss—to strengthen their understanding of real-world visual distributions and underlying physical dynamics.
Acknowledgments and Disclosure of Funding
We thank all the anonymous reviewers for their constructive comments. This work is supported in part by ByteDance Seed and National Natural Science Foundation of China (No. 62176002). Haoyuan Guo and Xu Sun are the corresponding authors of this paper.
Appendix A Impact Statement
@@
The evaluation of video generative models (VGMs) currently relies predominantly on human assessment or qualitative analysis. Our research demonstrates that multimodal large language models (MLLMs) could potentially serve as unified evaluators for AI-generated videos. The transition from human to automated evaluation frameworks would have significant dual implications: On the one hand, this shift would substantially reduce evaluation costs, thereby accelerating the development cycle of VGMs. On the other hand, if VGMs are optimized solely based on model feedback, there is a potential risk of divergence between the automated evaluators’ preferences and human values. Therefore, it essential to carefully scrutinize the correlation between automatic metrics and humans in the future research on AIGV evaluation.
Appendix B Related Work
B.1 Video Generative Models
The emergence of diffusion models [18, 53] has revolutionized generative applications, spanning from image synthesis [51, 48, 49] to video generation [16, 69, 19, 20]. On the one hand, recent VGMs like SORA [3], Kling [27], and MovieGen [46] have achieved remarkable success in producing highly realistic videos that closely resemble real ones. On the other hand, the condition signal for video generation has been expanded from text (T2V) to RGB images (I2V) [32, 6] and other videos (V2V) [78]. This work specifically focuses on evaluating text-conditioned video generation.
B.2 Multimodal Large Language Models
Multimodal large language models [45, 58, 29, 30] integrate LLMs with additional perception modules, enabling them to comprehend information beyond the text modality. With advancements in LLM backbones and the increasing scale of multimodal training, MLLMs are showing promising performance in understanding real-world images [36, 37, 13, 57] and videos [73, 31, 34]. However, understanding AI-generated visual content presents unique challenges compared to real images, and the application of MLLMs in this field is still in its early stages.
B.3 AI-Generated Video Evaluation
Despite the progress in VGMs, they still struggle to consistently generate high-quality videos, which necessitate a comprehensive and fine-grained evaluation. Initial evaluation approaches [21, 40] employ off-the-shelf models to evaluate specific aspects, such as CLIP [47] for video-text alignment and DINO [5] for frame-to-frame consistency. With the advancement of MLLMs, recent studies have incorporated these models in AIGV evaluation [17, 61, 56, 63, 2]. However, they are constrained to a fixed set of evaluation aspects and typically rely on human annotations to further fine-tune the MLLMs for AIGV evaluation. By contrast, in this work explores using MLLMs as unified evaluators for any AIGV aspect by leveraging their inherent vision-language understanding capabilities, eliminating the need for human annotations.
Appendix C More Details of Experimental Settings
C.1 Evaluated Systems
@@
We conduct zero-shot evaluations on 18 MLLMs, the detailed information of which is summarized in Tab. 7.
VideoScore-v1.1 [17].
This model is a fine-tuned version of Mantis-Idefics2-8B [23], trained to imitate human rating scores. In their work, He et al. [17] gathered human ratings (on a discrete scale from 1 to 4) for five key evaluation aspects: Visual Quality, Temporal Consistency, Dynamic Degree, Text-to-Video Alignment, and Factual Consistency. The VideoScore model is trained using a regression approach to predict these human rating scores. According to the definition of these aspects, we associate them with 8 out of the 15 aspects in UVE-Bench, as detailed in Tab. 8.
VBench [21].
VBench is a benchmark suite that comprehensively evaluates 16 aspects of AIGV using off-the-shelf models and specifically designed rules. As shown in Tab. 8, we identified an overlap of 6 aspects between VBench and our UVE-Bench, and we adopted the VBench metrics to evaluate these 6 aspects in UVE-Bench.
UMTScore [39].
UMTScore is designed to assess video–text alignment. It builds on the UMT-L/16 model [38], fine-tuned on the MSR-VTT dataset [64] for video–text retrieval. Within UVE-Bench, we employ UMTScore to evaluate the Overall Alignment aspect.
VIDEOCON-PHYSICS [2].
VIDEOCON-PHYSICS evaluates (1) whether a generated video adheres to physical commonsense and (2) whether it matches the given textual prompt. The model is initialized from VIDEOCON [1] and fine-tuned on a dataset of AI-generated videos with binary human annotations covering these two aspects. In UVE-Bench, we use VIDEOCON-PHYSICS to evaluate both Motion Naturalness and Overall Alignment.
DOVER [60].
DOVER assesses video quality from both technical and aesthetic perspectives. It is trained on DIVIDE-3k, a dataset of 3,590 real-world videos annotated with mean opinion scores (MOS). We adopt DOVER in UVE-Bench to evaluate Technical Quality and Aesthetic Quality.
C.2 Evaluation Criteria for Single Video Rating
@@
[…]
@@
where $s$ is the coefficient that controls the speed of decaying. The graph of $f_{c}(\mathcal{S}|\beta)$ and $f^{\prime}_{c}(\mathcal{S}|\alpha)$ are illustrated in Fig. 4. An analysis of the impact of $\alpha$ and $\beta$ on the final result is conducted in Appendix D.5.
C.3 Methodology for Adapting Single Video Rating for Pairwise Comparison
@@
In the above equations, $\alpha<\beta$ are the thresholds for “bad” and “good” videos, respectively. $\tau$ is a threshold of the difference between $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, which controls whether we should select from $\{``\mathcal{V}_{1}\text{better}",``\mathcal{V}_{2}\text{better}$"\} or $\{``\text{same good}",``\text{same bad}$"\}. By default, we set $\alpha=0.4,\beta=0.8,\tau=0.05$.
C.4 Annotation Interface
Fig. 5 presents the annotation interface used for establishing our human baseline. Annotators are shown a pair of videos along with aspect-specific evaluation instructions. Based on these, they indicate their preference by selecting one of four options: {“A is better”, “B is better”, “same good”, “same bad”}.
Appendix D More Experimental Results
D.1 Effect of Prompting Strategy
@@
Fig. 11(a) presents the results of different prompting strategy with InternVL-2.5-8B-MPO, Qwen2-VL-7B and Qwen2-VL-72B. Unlike InternVL-2.5-78B-MPO (previously presented in Fig. 4(b)), these three models showed a less pronounced advantage when using detailed aspect-specific descriptions, with the simplified prompt achieving better performance in several aspects. We conjecture that this is because these three models cannot comprehend the detailed description as effectively as InternVL-2.5-78B-MPO. Nevertheless, when considering overall performance, the full prompt consistently maintained a slight improvement over the simplified version across all three models. Consistent with our findings for InternVL-2.5-78B-MPO, removing video order indicator does not result in significant performance change.
D.2 Effect of Scoring Strategy
@@
In §4.3.2, we analyze the effect of different scoring strategies with Qwen2-VL-7B. Here, we extended the investigation to three additional models: Qwen2-VL-72B, InternVL-2.5-8B-MPO, and InternVL-2.5-78B. The results are presented in Fig. 11(b). Our findings reveal consistent patterns across all models. Similar to Qwen2-VL-7B, the good/bad and yes/no scoring strategies demonstrate distinct strengths in different evaluation aspects. Direct score generation through MLLMs consistently yields suboptimal results across all models. However, compared with Qwen2-VL-7B, the advantage of adaptive scoring token is less pronounced in Qwen2-VL-72B, InternVL-2.5-8B-MPO and InternVL-2.5-78B. This suggests that these three models are more robust to the variation of scoring tokens. These findings confirm that using yes/no as unified scoring token is simple yet effective.
D.3 Negative Prompt versus Positive Prompt
By default, we prompt MLLMs to assess whether a video is “good” in a specific aspect. As an alternative, we can also frame the question in a negative form—asking whether a video is “bad” in that aspect. To examine the impact of this change, we modify the evaluation prompts accordingly and invert the positive and negative token assignments when computing single video ratings in Eq. 1.
We then evaluate the Qwen2-VL-7B model using both positive and negative prompts across four aspects. As shown in Table 11, the results reveal clear differences across dimensions. For Static Quality and Temporal Quality, the performance gap between positive and negative prompts is relatively small. In contrast, for Dynamic Degree and Video-Text Alignment, positive prompts yield substantially higher performance. These findings suggest that positive prompts provide more stable and reliable evaluations. Therefore, we adopt positive prompts as the default configuration in our experiments.
D.4 Adapting Single Ratings for Pairwise Comparison
While our previous experiments involve direct pairwise video comparisons by simultaneously feeding two videos into MLLMs, we now explore an alternative approach. This method first independently rates each video using Eq. 1 and then converts these individual ratings into a four-way selection from set $\mathbf{O}$. The detailed conversion methodology is discussed in Appendix C.3.
As shown in Fig. 4(d), adaptation from single video rating brings substantial improvement to Qwen2-VL-7B compared to directly performing video pair comparison. However, Qwen2-VL-72B shows only minimal improvements under the same approach. Similar patterns were observed with InternVL-2.5-MPO, as shown in Appendix D.4. We hypothesize that smaller-scale models, like the 7B variant, have inherently weaker pairwise comparison capabilities, and adapting from single video rating bypasses this limitation. Based on these findings, we recommend implementing the single-to-pairwise adaptation strategy particularly for MLLMs with limited pairwise comparison capabilities.